OpenStudy (anonymous):

Y^2 = X^3 - 432 12^2 = 36^3 - 432 is?

OpenStudy (cgreenwade2000):

If you guessed those for x and y, you may want to try again.

OpenStudy (anonymous):

Explain, I guessed, correct. But im confused on how to even go about solving it. Y^2 = X^3 - 432.

OpenStudy (anonymous):

But, Y^2 is 144 ( with Y = 12 ) 432 - 144 = ?, ? would equal X^3?

hartnn (hartnn):

432-144 = 288 12^2+144=24^2

OpenStudy (anonymous):

Your saying that 24^2 Is X^3? Wouldnt that be making the power incorrect?

hartnn (hartnn):

do u need to find X ?

OpenStudy (anonymous):

You want to find Y^2, Y^2 = X^3 - 432. So, yes. Forget about Y^2, once we find X^3 we will have Y^2.

hartnn (hartnn):

aren't u given anything else ?

OpenStudy (anonymous):

No, I was not. Just Y^2 = X^3 - 432. But assuming Y = 12. 12^2 = 144. 144- 432 = 288. What to the power of 3 equals 288? :/

hartnn (hartnn):

nothing actually...so your assumption that y=12 is not correct

OpenStudy (anonymous):

ok... so we need to find a new Y^2... What can be powered by 3 and still be under 432, and whats left over still be powered by 2?

hartnn (hartnn):

12^3-432 = 1296 sqrt{1296} = 36 Y=12 X=36

OpenStudy (anonymous):

I dont understand your thoughts. Explain,

hartnn (hartnn):

sorry, X=12 Y=36

hartnn (hartnn):

i tried different values of X, from 7 onwards, when i reached 12, 1728-432 = 1296 which was perfect square

OpenStudy (anonymous):

So your saying I had it backwards earlier? ^^^ So, How would that be put into equation? Y^2 = X^3 - 432. 12^3 - 432 = 1296 SR - ( 1296 ) = 36 X= 12 Y= 36. Thats it? Thats the answer?

hartnn (hartnn):

yup...

OpenStudy (anonymous):

Whered you get 12 from?